The product of two consecutive integers is 5 less than 5 times their sum. Find the integers
Answer by lwsshak3(11628)

The product of two consecutive integers is 5 less than 5 times their sum. Find the integers
**
let x=first digit
x+1=consecutive digit
..
x(x+1)=5(x+x+1)-5
x^2+x=10x+5-5
x^2-9x=0
x(x-9)=0
x=0 (reject, not a solution)
or
x=9
x+1=10
first digit=9
consecutive digit=10